1. Suppose that E and F are mutually exclusive events, Pr(E) = 0.4, Pr(E) = 0.5

what is Pr(E U F)?

2. Your drawer contains 8 red socks and 6 green socks. it is too dark to see which are which.

Choose 2 socks. What is the probability that you pick a pair of socks ( either 2 red or 2 green)

3. In a school of 1200 students, 250 are seniors, 150 students take math, and 40 students are seniors and are also taking math

what is the probability that a randomly chosen student who is a senior, is taking math?

4. Cross-fertilizing a red and a white flower produces red flowers 25% of the time. now we cross-fertilize pairs of red

and white flowers and produce five offspring. what is the probability that there are no red flower plants in the five offspring?
Answer #1

1. Since E and F are two mutually exclusive events, En F = 0 P(ENF) = 0 We know; PEUF) = P(E) + P(F) - PEF = P(E) + P(F) sinc

2. There are 8 red socks and 6 green socks, thus total of 14 socks Number of ways in which one can draw a pair of socks (14)

3. In a school there are 1200 students, 250 seniors, 150 taking math and 40 students seniors and taking maths Probability of

4. Probability of producing a red flower due to cross fertilizing a red and white flower is 0.25

Probability of getting no red flower in the five offsprings from cross fertilization is

(0.25)5 = 0.00098